# Antidote: Post-fine-tuning Safety Alignment for Large Language Models against Harmful Fine-tuning

### v2

@@ Abstract
Safety aligned Large Language Models (LLMs) are vulnerable to harmful fine-tuning attacks (Qi et al. 2023)– a few harmful data mixed in the fine-tuning dataset can break the LLMs’s safety alignment. Existing mitigation strategies include alignment stage solutions (Huang, Hu, and Liu 2024; Rosati et al. 2024a) and fine-tuning stage solutions (Huang et al. 2024; Mukhoti et al. 2023). However, our evaluation shows that both categories of defenses fail when some specific training hyper-parameters are chosen – a large learning rate or a large number of training epochs in the fine-tuning stage can easily invalidate the defense, which however, is necessary to guarantee finetune performance. To this end, we propose Antidote, a post-fine-tuning stage solution, which remains agnostic to the training hyper-parameters in the fine-tuning stage. Antidote relies on the philosophy that by removing the harmful parameters, the harmful model can be recovered from the harmful behaviors, regardless of how those harmful parameters are formed in the fine-tuning stage. With this philosophy, we introduce a one-shot pruning stage after harmful fine-tuning to remove the harmful weights that are responsible for the generation of harmful content. Despite its embarrassing simplicity, empirical results show that Antidote can reduce harmful score while maintaining accuracy on downstream tasks. Our project page is at https://huangtiansheng.github.io/Antidote_gh_page/
Introduction
@@
Fine-tuning-as-a-service has become a new paradigm for Large Language Model (LLM) service with an increasing demand for personalized service delivery. Typically, fine-tuning data are uploaded by the users, and the service provider (e.g., OpenAI) finetunes a pre-trained LLM (foundation model), which is then served to meet the users’ customized need.
@@
Before fine-tuning for user tasks, a pre-trained LLM is usually safety aligned to guarantee that the outputs of the LLM meet the safety preference, i.e., to refuse to generate harmful content even when the users trigger them to do so. However, recent studies (Qi et al. 2023; Yang et al. 2023; Zhan et al. 2023) show that a few harmful data mixed in the fine-tuning dataset can trigger the model to forget the alignment knowledge it learned previously –it no longer uses refusal response when users submit a harmful prompt.
@@
Figure 1: Antidote with a three-stage pipeline, i.e., i) safety alignment, ii) user fine-tuning, iii) one-shot pruning. While existing works mostly focus on optimization in the first stage, e.g., (Huang, Hu, and Liu 2024; Rosati et al. 2024a) or the second stage (Huang et al. 2024; Mukhoti et al. 2023), Antidote utilizes the post-fine-tuning stage to prune the harmful weights to recover the model from harmful behaviors.
@@
Existing mitigation strategies can be broadly categorized into two categories, i.e., alignment stage defense and user fine-tuning stage defense. The first category is concerned with how to improve the large language model’s immunization towards the harmful fine-tuning data in the alignment stage. For example, (Huang, Hu, and Liu 2024) add artificial perturbation in the alignment stage to simulate the harmful embedding drift in the fine-tuning stage, and utilizes a mini-max optimization to enforce the model to be immune to the perturbation. (Rosati et al. 2024b) utilize a representation noising technique to degrade the representation distribution of the harmful data to a random Gaussian noise, such that the harmful content generation is more difficult to learn by harmful fine-tuning data. For fine-tuning-stage mitigation, the core idea is to mitigate the forgetting of the alignment knowledge but also to learn the knowledge for the users’ tasks. To achieve this goal, (Huang et al. 2024) separate the fine-tuning stage into two states, which alternatively optimize the alignment and fine-tuning dataset. (Mukhoti et al. 2023) add a regularization to constrain drift in feature space to mitigate the forgetting of alignment knowledge. However, our empirical evaluation in Section III reveals a common weakness of the existing defense methods – a small learning rate and a small number of epochs in the fine-tuning stage are required to guarantee their effectiveness. This requirement can be detrimental to downstream tasks’s performance because some fine-tuning tasks require a larger learning rate and longer training epochs to guarantee learning performance.
To this end, we in this paper aim to answer the following research question:
@@
Is there a safety alignment that can be insensitive to the hyper-parameters for the fine-tuning stage?
@@
Guiding by this urge, we propose Antidote, a safety alignment solution that realigns the model after the fine-tuning stage has been fully completed. The design of Antidote is agnostic to how the fine-tuning is done – it relies on the philosophy that by removing the harmful parameters, the harmful model can be recovered from the harmful behaviors, regardless of how those harmful parameters are formed in the fine-tuning stage. Empirically, we show that Antidote respectively reduces the harmful score by up-to 17.8% (compared to SFT without defense) while maintaining the same level of fine-tuning accuracy (by up-to 1.83% accuracy loss).
@@
Safety alignment. Safety alignment is about how to align a large language model such that its outputs are aligned with humans’ values. Representative techniques are RLHF (Ouyang et al. 2022) and its variants (Dai et al. 2023; Bai et al. 2022; Wu et al. 2023; Dong et al. 2023; Rafailov et al. 2023; Yuan et al. 2023). Most recently, there are alternative solutions focusing on augmenting the alignment data, e.g., (Liu, Sferrazza, and Abbeel 2023; Liu et al. 2023; Ye et al. 2023).
@@
Harmful fine-tuning. (Qi et al. 2023; Yang et al. 2023; Zhan et al. 2023) show that LLMs aligned by RLHF or SFT (supervised fine-tuning) can be jail-broken after fine-tuning on explicit/implicit harmful data, and several mechanism studies (Leong et al. 2024; Wei et al. 2024; Peng et al. 2024; Jain et al. 2024) are conducted to analyze the problem. Existing solutions for harmful fine-tuning issues can be categorized into two main categories. The first category is alignment stage solutions, which study how to improve the model’s immunization ability to the fine-tuning by modifying training procedure in alignment stage. Exemplar literature are Vaccine (Huang, Hu, and Liu 2024) and RepNoise (Rosati et al. 2024b, a). Specifically, Vaccine vaccinates the model by adding embedding perturbation in the alignment stage, and RepNoise improve the robustness by enforcing the representation of the harmful data to be a random Gaussian noise.A concurrent research (Tamirisa et al. 2024) utilize meta learning to immunize the model. The second category is fine-tuning-stage solutions (Mukhoti et al. 2023; Zong et al. 2024; Huang et al. 2024; Wang et al. 2024; Lyu et al. 2024)., which study how to avoid forgetting the alignment knowledge while also learning the fine-tuning knowledge by modifying training procedure in user fine-tuning stage. Specifically, LDIFS (Mukhoti et al. 2023) introduces a regularizer to enforce the iterate’s embedding to be in close proximity to that of the aligned model. VLGuard(Zong et al. 2024) mixes the alignment data with the user data to train in the user fine-tuning stage. Lisa(Huang et al. 2024) alternatively optimizes over the alignment data and the user fine-tuning data and use proximal regularizer to enforce proximity between iterates. Recently, there are also more advanced attack (He, Xia, and Henderson 2024; Halawi et al. 2024) in this setting.
@@
Model sparsification. Since (Frankle and Carbin 2018), model sparsification for deep learning models has been extensively studied. The core research problem for model sparsity literature is to score the weights coordinates according to their importance. The coordinates with small importance score are then removed to compress the model. Specifically for large language model, (Frantar and Alistarh 2023) propose SparseGPT, which forms the importance score by solving a layer-wise reconstruction problem. (Sun et al. 2023) propose Wanda score, which utilize joint weights/activation metrics to measure the coordinate importance. On top of Wanda, (Yin et al. 2023) propose layer-wise sparsity, which further improve the model compression ratio. We in this paper borrow importance score from the model sparsification literature to identify and remove harmful parameters.
@@
We acknowledge that there are a few concurrent post-fine-tuning stage defenses, which also aim at purifying the model after fine-tuning completes. For example, (Hsu et al. 2024) project the harmful gradient updates to a aligned subspace. (Yi et al. 2024) realign model via subspace-oriented model fusion. (Tong et al. 2024) realign by self-contrastive decoding. It is possible that these existing solutions can also be insensitive the training hyper-parameters in fine-tuning stage due to the good nature of post-fine-tuning solutions. However, prior to this study, there is no systematical study on the hyper-parameter sensitivity issue, which highlights the significance of post-fine-tuning stage defense.
Preliminaries
Threat Model and Assumptions
@@
A typical fine-tuning-as-a-service pipeline. is considered in (Qi et al. 2023), where the users upload a few fine-tuning data to the service provider. On behalf of users, the service provider fine-tunes the aligned pre-trained model on this dataset, and the finetuned model is deployed by the service provider to deliver personalized service. Because the model is deployed in the service provider’s server and the answer to user prompts is delivered by the service provider’s API, the service provider has an obligation to ensure the answer is harmless. Otherwise, the service provider might face governance issues (Reuel et al. 2024) or lawsuit  [^1] Laws have been enforced, e.g., SB-1047 in California..
@@
Following (Rosati et al. 2024a; Tamirisa et al. 2024), we assume the service provider hosts a harmful dataset $\mathcal{D}_{realign}$ (containing harmful prompt-harmful answer pairs), which we use to perform post-fine-tuning stage re-alignment. We henceforth refer to this dataset as re-alignment dataset for clearness. Following (Rosati et al. 2024a; Huang, Hu, and Liu 2024; Hsu et al. 2024; Zong et al. 2024), we assume the service provider maintains a safety alignment dataset $\mathcal{D}_{align}$ (containing harmful prompt-safe answer pairs).
Revisiting Existing Defenses
@@
In this subsection, we evaluate the existing safety alignment for harmful fine-tuning issue. We choose two representative alignment-stage solutions (Huang, Hu, and Liu 2024; Rosati et al. 2024a) and two fine-tuning-stage solutions (Huang et al. 2024; Mukhoti et al. 2023) as demonstration.
Figure 2: Harmful score and finetune accuracy with different learning rates in the fine-tuning stage. Here we fix fine-tuning epochs to 20. Harmful score and finetune accuracy with different learning rates in the fine-tuning stage. Here we fix fine-tuning epochs to 20.
@@
Existing defenses fail with a large learning rate in fine-tuning stage. We adjust the learning rate in the user fine-tuning stage and show how existing methods perform in Figure 2. Our results show that both alignment-stage defenses (Vaccine and RepNoise) and fine-tuning-stage defenses (Lisa and LDIFS) tend to have larger harmful scores when learning rate is large. We now explain the reason for their failures. i) For alignment stage solutions, the core idea of defense is to strengthen the aligned model’s robustness towards harmful data in the later fine-tuning stage. The reason for a degraded performance is that a larger learning rate in fine-tuning stage can make it easier to subvert the model robustness induced before fine-tuning (the same phenomenon and explanation are given in (Rosati et al. 2024a)). ii) For fine-tuning stage defenses, the core idea is to introduce a regularizer in the fine-tuning stage to enforce the fine-tuning iterate in proximity to the aligned model. These solutions also suffer degraded performance because a larger learning rate may drift the iterates far away from the aligned model, resulting in the model failing to converge near the aligned model.
@@
Figure 3: Harmful score and finetune accuracy with different fine-tuning epochs in the user fine-tuning stage. Here we fix fine-tuning learning rate to 1e-5.
@@
A sufficiently large learning rate and finetune epochs are necessary. However, as shown in Table 2 and 4, a sufficiently large learning rate is necessary to guarantee good fine-tune accuracy, which indicates that state-of-the-art solutions fall short and need renovation.
@@
Methodology
In order to counter the hyper-parameters sensitivity issue in fine-tuning stage, we propose a post-fine-tuning safety alignment that remains agnostic to the exact training setting in fine-tuning. The high-level idea is to remove the harmful parameters in the model after the model has corrupted with fine-tuning. The method is agnostic to the hyper-parameter in fine-tuning stage because in principle the harmful parameters can anyway be deactivated regardless of how they form in the fine-tuning stage. We refer to Figure 4 for a system overview.

### v3

@@ Abstract
Safety aligned Large Language Models (LLMs) are vulnerable to harmful fine-tuning attacks – a few harmful data mixed in the fine-tuning dataset can break the LLMs’s safety alignment. While several defenses have been proposed, our evaluation shows that existing defenses fail when some specific training hyper-parameters are chosen – a large learning rate or a large number of training epochs in the fine-tuning stage can easily invalidate the defense. To this end, we propose Antidote, a post-fine-tuning stage solution, which remains agnostic to the training hyper-parameters in the fine-tuning stage. Antidote relies on the philosophy that by removing the harmful parameters, the harmful model can be recovered from the harmful behaviors, regardless of how those harmful parameters are formed in the fine-tuning stage. With this philosophy, we introduce a one-shot pruning stage after harmful fine-tuning to remove the harmful weights that are responsible for the generation of harmful content. Despite its embarrassing simplicity, empirical results show that Antidote can reduce harmful score while maintaining accuracy on downstream tasks. Code is available at https://github.com/git-disl/Antidote.
Machine Learning, ICML
1 Introduction
@@
Fine-tuning-as-a-service has become a new paradigm for Large Language Models (LLM) service with an increasing demand for personalized service delivery. Typically, fine-tuning data are uploaded by the users, and the service provider (e.g., OpenAI) finetunes a pre-trained LLM, which is then served to meet the users’ customized need.
@@
Before fine-tuning for user tasks, a pre-trained LLM is usually safety aligned to guarantee that the outputs of the LLM meet the safety preference, i.e., to refuse to generate harmful content even when the users trigger them to do so. However, recent studies (Qi et al., 2023; Yang et al., 2023; Zhan et al., 2023; Lermen et al., 2023; Yi et al., 2024a) show that a few harmful data mixed in the fine-tuning dataset can trigger the model to forget the alignment knowledge it learned previously –it no longer uses refusal response when users submit a harmful prompt.
@@
Figure 1: Antidote with a three-stage pipeline, i.e., i) safety alignment, ii) user fine-tuning, iii) one-shot pruning. While existing defenses focus on the first stage, e.g., (Huang et al., 2024d; Rosati et al., 2024a) or the second stage (Huang et al., 2024b; Mukhoti et al., 2023), Antidote utilizes the post-fine-tuning stage to prune the harmful weights to recover the model from harmful behaviors.
@@
Existing mitigation strategies can be broadly categorized into two categories, i.e., alignment stage defense and user fine-tuning stage defense. The first category is concerned with how to improve the large language model’s immunization towards the harmful fine-tuning data in the alignment stage. For example, (Huang et al., 2024d) add artificial perturbation in the alignment stage to simulate the harmful embedding drift in the fine-tuning stage, and utilizes a mini-max optimization to enforce the model to be immune to the perturbation. (Rosati et al., 2024b) utilize a representation noising technique to degrade the representation distribution of the harmful data to a random Gaussian noise, such that the harmful content generation is more difficult to learn by harmful fine-tuning data. For fine-tuning-stage mitigation, the core idea is to mitigate the forgetting of the alignment knowledge but also to learn the knowledge for the users’ tasks. To achieve this goal, (Mukhoti et al., 2023) add a regularization to constrain drift in feature space to mitigate the forgetting of alignment knowledge and (Huang et al., 2024b) separate the fine-tuning stage into two states, which alternatively optimize the alignment and fine-tuning dataset.
However, our empirical evaluation in Section III reveals a common weakness of the existing defense methods – a small learning rate and a small number of epochs in the fine-tuning stage are required to guarantee their effectiveness. This requirement can be detrimental to downstream tasks’s performance because some fine-tuning tasks require a larger learning rate and longer training epochs to guarantee learning performance. To this end, we in this paper aim to answer the following research question:
@@
Is there a defense that can be less sensitive to the hyper-parameters of the fine-tuning stage?
@@
Driven by this question, we propose Antidote, a defense that realigns the model after the fine-tuning stage has been fully completed. The design of Antidote is agnostic to how the fine-tuning is done – it relies on the philosophy that by removing the harmful parameters, the harmful model can be recovered from the harmful behaviors, regardless of how those harmful parameters are formed in the fine-tuning stage. Empirically, we show that Antidote respectively reduces the harmful score by up-to 17.8% (compared to SFT without defense) while maintaining the same level of fine-tuning accuracy (by up-to 1.83% accuracy loss).
@@ 2 Related Work
Safety alignment. Safety alignment is about how to align an LLM such that its outputs are aligned with humans’ values. Representative techniques are RLHF (Ouyang et al., 2022) and its variants (Dai et al., 2023; Bai et al., 2022; Wu et al., 2023; Dong et al., 2023; Rafailov et al., 2023; Yuan et al., 2023). Most recently, there are alternative solutions focusing on augmenting the alignment data, e.g., (Liu et al., 2023a, b; Ye et al., 2023; Tekin et al., 2024).
@@
Harmful fine-tuning. (Qi et al., 2023; Yang et al., 2023; Zhan et al., 2023; Lermen et al., 2023; Yi et al., 2024a) show that LLMs aligned by RLHF or SFT (supervised fine-tuning) can be jail-broken after fine-tuning on explicit/implicit harmful data, and several mechanism studies (Leong et al., 2024; Wei et al., 2024; Peng et al., 2024; Jain et al., 2024; Qi et al., 2024b; Hsiung et al., 2025; Guo et al., 2024; Poppi et al., 2024; Che et al., 2025; Chen et al., 2025) are conducted to analyze the problem. Existing solutions for harmful fine-tuning can be categorized into two categories. The first category is alignment stage solutions, which study how to improve the model’s immunization ability to the fine-tuning by modifying training procedure in alignment stage. Examples are Vaccine (Huang et al., 2024d) and RepNoise (Rosati et al., 2024b, a). Vaccine vaccinates the model by adding embedding perturbation in the alignment stage, and RepNoise improved the robustness by enforcing the representation of the harmful data to be a random Gaussian noise. Other alignment solutions include CTRL (Liu et al., 2024c), TAR (Tamirisa et al., 2024), Booster (Huang et al., 2024a), SN-Tune (Zhao et al., 2025b), T-Vaccine (Liu et al., 2024a), CTRAP (Yi et al., 2025b), KT-IPA (Cheng et al., 2025), SAM unlearning (Fan et al., 2025), Reward Neutralization (Cao, 2025) and SEAM (Wang et al., 2025c). The second category is fine-tuning-stage solutions (Mukhoti et al., 2023; Bianchi et al., 2023; Zong et al., 2024; Huang et al., 2024b; Wang et al., 2024; Lyu et al., 2024; Qi et al., 2024a; Shen et al., 2024; Choi et al., 2024; Du et al., 2024; Li et al., 2025; Eiras et al., 2024; Li & Kim, 2025; Li et al., 2024b; Liu et al., 2024b; Zhao et al., 2025a; Liu et al., 2025; Li, 2025; Wu et al., 2025; Peng et al., 2025), which study how to avoid forgetting the alignment knowledge while also learning the fine-tuning knowledge by modifying training procedure in user fine-tuning stage. Specifically, LDIFS (Mukhoti et al., 2023) introduces a regularizer to enforce the iterate’s embedding to be in close proximity to that of the aligned model. Lisa (Huang et al., 2024b) alternatively optimizes over the alignment data and the fine-tuning data and use proximal regularizer to enforce proximity between iterates. Recently, there are advanced attacks (He et al., 2024; Halawi et al., 2024; Guan et al., 2025; Huang et al., 2025a; Davies et al., 2025; Kazdan et al., 2025), and there are attacks towards other settings, e.g., federated learning(Ye et al., 2024; Li et al., 2024a), diffusion models (Pan et al., ) and large reasoning model (Huang et al., 2025b). For a more comprehensive discussion, we refer to surveys (Huang et al., 2024c; Wang et al., 2025a; Verma et al., 2024)
@@
Model sparsification. Since (Frankle & Carbin, 2018), model sparsification for deep learning models has been extensively studied. The core research problem for model sparsity is to score the weights coordinates according to their importance, and then remove the unimportant ones to compress the model. For LLMs, (Frantar & Alistarh, 2023) propose SparseGPT, which forms the importance score by solving a layer-wise reconstruction problem. (Sun et al., 2023) propose Wanda score, which utilize joint weights/activation metrics to measure the coordinate importance. On top of Wanda, (Yin et al., 2023) propose layer-wise sparsity, which further improve the model compression ratio. We in this paper borrow importance score from the model sparsification literature to identify and remove harmful parameters.
@@
We acknowledge that there are a few concurrent post-fine-tuning stage defenses, aiming at purifying the model after fine-tuning completes. RESTA (Bhardwaj et al., 2024) realigns the model by interpolating a safety vector to the compromised model. LAT (Casper et al., 2024) utilize embedding space perturbation to unlearn the harmful knowledge, Safe LoRA (Hsu et al., 2024) projects the harmful updates to an aligned subspace. SOMF (Yi et al., 2024c) realigns model via subspace-oriented model fusion. (Tong et al., 2024) realign by self-contrastive decoding. IRR (Wu et al., 2024) and NLSR (Yi et al., 2024b) realigns by neuron correction, SafetyLock (Zhu et al., 2024) realigns by activation patching, and Panacea (Wang et al., 2025b) optimizes the post-fine-tuning perturbation that maximally increases the safety loss. There are several other post-fine-tuning stage solutions that are worth to be checked out, e.g., (Yi et al., 2025a; Liu et al., 2024b; Wu et al., 2024; Gong et al., 2025; Djuhera et al., 2025; Yang et al., 2025; Lu et al., 2025). It is possible that these concurrent defense can also be insensitive the hyper-parameters in fine-tuning stage. However, prior to us, there is no systematical study on hyper-parameter sensitivity issue, which highlights the significance of post-fine-tuning stage defense.
3 Preliminaries
3.1 Threat Model and Assumptions
@@
Fine-tuning-as-a-service. Fine-tuning-as-a-service is illustrated in Figure 1. In this scenario, users upload fine-tuning data to the service provider. On behalf of users, the provider fine-tunes the aligned pre-trained model on this dataset, and the finetuned model is deployed to deliver personalized service to users. The fine-tuned data is uploaded by users and therefore incurring safeyt risk. Because the model is deployed in the service provider’s server and the answer to user prompts is delivered by the service provider’s API, the service provider has an obligation to ensure the answer is harmless. Otherwise, the provider might face governance issues (Reuel et al., 2024; Huang et al., 2024c) or lawsuit  [^1] Regulations, e.g., SB-1047 in California, are considered..
@@
Assumptions. We assume the service provider hosts a harmful dataset $\mathcal{D}_{realign}$ (containing harmful prompt-harmful answer pairs), which we use to perform post-fine-tuning stage re-alignment. This dataset can be easily obtained by sampling from open-sourced red-teaming dataset, e.g., BeaverTails (Ji et al., 2023), HH-RLHF, etc. Of note, such a harmful dataset is also assumed in already accepted papers (Rosati et al., 2024a; Huang et al., 2024a) and a prior work (Tamirisa et al., 2024), and therefore should not be too strong or out of generality. We henceforth refer to this dataset as re-alignment dataset for clearness. Following (Rosati et al., 2024a; Huang et al., 2024d; Hsu et al., 2024; Zong et al., 2024), we assume the service provider maintains a safety alignment dataset $\mathcal{D}_{align}$ (containing harmful prompt-safe answer pairs).
3.2 Hyper-parameter Sensitivity Issue
@@
In this subsection, we evaluate the existing safety alignment for harmful fine-tuning issue and identify their insufficiency. We choose two representative alignment-stage solutions (Huang et al., 2024d; Rosati et al., 2024a) and two fine-tuning-stage solutions (Huang et al., 2024b; Mukhoti et al., 2023) as demonstration.
@@
Existing defenses fail with a large learning rate in fine-tuning stage. We adjust the learning rate in the user fine-tuning stage and show how existing methods perform in Figure 2. Our results show that both alignment-stage defenses (Vaccine and RepNoise) and fine-tuning-stage defenses (Lisa and LDIFS) tend to have larger harmful scores when learning rate is large. We now explain the reason for their failures. i) For alignment stage solutions, the core idea of defense is to strengthen the aligned model’s robustness towards harmful data in the later fine-tuning stage. The reason for a degraded performance is that a larger learning rate in fine-tuning stage can make it easier to subvert the model’s safety alignment (the same phenomenon and explanation are given in (Rosati et al., 2024a)). ii) For fine-tuning stage defenses, the core idea is to introduce a regularizer in the fine-tuning stage to enforce the fine-tuning iterate in proximity to the aligned model. These solutions also suffer degraded performance because a larger learning rate may drift the iterates far away from the aligned model, resulting in the model failing to converge near the aligned model.
@@
Figure 2: Harmful score and finetune accuracy with different learning rates after fine-tuning. Here we fix fine-tuning epochs to 20.
Figure 3: Harmful score and finetune accuracy with different fine-tuning epochs after user fine-tuning. Here we fix fine-tuning learning rate to 1e-5.
@@
A sufficiently large learning rate and finetune epochs are necessary. However, as shown in the right figures of Table 2 and 4, a sufficiently large learning rate is necessary to guarantee good fine-tune accuracy, which indicates that state-of-the-art solutions fall short and need renovation.
@@
4 Methodology
In order to counter the hyper-parameters sensitivity issue in fine-tuning stage, we propose a post-fine-tuning safety alignment that remains agnostic to the exact training setting in fine-tuning.
The high-level idea of the proposed defense, named Antidote, is to remove the harmful parameters in the model after the model has been corrupted with fine-tuning. The method is agnostic to the hyper-parameter in fine-tuning stage because in principle the harmful parameters can anyway be deactivated regardless of how they form in the fine-tuning stage. We refer to Figure 4 for a system overview.